# Leveraging Analog Neutral Atom Quantum Computers for Diversified Pricing in Hybrid Column Generation Frameworks

### v1

@@ IV.2 Pulse Design
We benchmark the performance of these two pulse strategies in terms of quality and diversity across 20 different MWIS instances of weighted Erdos-Rényi random graphs with $|V|=30$ nodes and edge probability $p$. The weights of the nodes were drawn uniformly from the interval $[1,10]$. Manipulation of the quantum pulses is done using the Pulser [56] Python packages. Emulations of $N\leq 12$ qubits are done using the QuTip backend [57], while those with $N>12$ are done using emu-mps [58], an open-access tool for matrix-product state (MPS) simulations of neutral atom quantum systems. The emulator’s parameters are kept in their default settings.
@@ V.1 Synthetic instances
At each iteration of the CG scheme, there are $|V|$ PSPs that must be solved over subgraphs $G_{v}$ for all $v\in V$. To ensure a control over the hardness of our instances, we make sure that all $G_{v}$ remain hard (i.e. non-trivial).
We control the number of vertices $\lvert K_{v}\rvert$ (i.e. the number of compatible tours per vehicle class), the number of edges $\lvert E_{v}\rvert$ via the edge’s probability $p=\lvert E_{v}\rvert/\binom{\lvert K_{v}\rvert}{2}$ (i.e. the number of conflicting tours).
We also control the number of vehicle classes $\lvert V\rvert$ (i.e. the number of independent PSPv to solve at each iteration). There are also parameters influencing the complexity of the RMP: the average number of allowed vehicle types per tours $\bar{N_{V}}=1/\lvert K\rvert\sum_{k=1}^{\lvert K\rvert}|V_{k}|$, from which we infer the total number of tours in the problem as $K=\lceil\bar{N_{V}}\lvert K\rvert\rceil$, the costs of each tour $C_{l}^{(T)}$ and vehicle types $C_{v}^{(v)}$, and the availability bounds $N_{v}^{\min}$ and $N_{v}^{\max}$ for each vehicle type $v$. Our synthetic instances have the following global parameters: $\lvert V\rvert=8$, graphs $G_{v}$ are Erdos-Renyi connected graphs with $p\sim 0.30$, $\bar{N_{V}}=2$, $C_{k}^{(T)}\in\mathcal{N}(10,5)$, $C_{v}^{(v)}\in\mathcal{N}(50,10)$, $N_{v}^{\min}=1$ and $N_{v}^{\max}$ where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and variance $\sigma^{2}$. The Instance classes are noted by the tuple $(|V|,|K_{v}|)$. We create 10 random instance (random generation of subgraphs and costs following the above distributions), for the $(8,8)$, $(8,10)$ and $(8,12)$ instances. For the $(8,14)$ instances, we generate 30 instances in order to perform a more in-depth analysis of the impact of the chosen solver’s performance on the CG scheme.
@@ VI Conclusion
We showed in this paper that novel register embedding and pulse design protocols for analog neutral atom quantum computers can enhance the performance of hybrid column generation schemes. We showed that these protocols can compete with other classical solvers for the pricing subproblems in returning a large set of high quality and diversity columns at each iteration. We also introduced a new greedy postprocessing method called Make_Diff which seeks to modify degenerate samples in order to obtain a set of potential columns that are all non-degenerate. Importantly, our quantum protocol with this postprocessing most closely mimics the behavior of an idealized solver that finds the $M$ best columns to add to the reduced master problem.
The focus of this paper was on the fleet assignment problem, which acted as a hard combinatorial problem on which to test our hypotheses concerning the design and performance of analog quantum primitives in column generations schemes. We note that the conclusions and the methods presented here can be applied to other practical settings. In machine job scheduling, for example, jobs must be assigned to machines in a way that respects machine capabilities and job durations, aiming to optimize various cost metrics. Likewise, in airline and public transit scheduling, aircraft, or buses must be assigned to flight or route sequences while accounting for various constraints.
In such industrial settings, one may need to expand the workflow to include branching (as was done in Ref. [60]) or the generation of cuts [61].
These settings all represent prime places on which an industrial quantum advantage may be sought. To achieve such an advantage, our paper places a threshold to meet: the quantum protocol must be able to return a set of $M$ high quality and high diversity columns at each iterations faster than competing classical methods, such as simulated annealing and tensor networks.
This remains an open question, although the experimental results of quantum annealing protocols to solve large-scale maximum independent set problems on neutral atom hardware [62, 55, 15] are encouraging.
@@ Acknowledgements.
We thank Wesley Coelho, Yassine Naghmouchi and Hossein Sadeghi for valuable discussions and support. This research was financially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and Pasqal Canada through an Alliance grant. This work made use of compute resources by Calcul Québec and the Digital Research Alliance of Canada.

### v2

@@ IV.2 Pulse Design
We benchmark the performance of these two pulse strategies in terms of quality and diversity across 20 different MWIS instances of weighted Erdos-Rényi random graphs with $|V|=30$ nodes and edge probability $p$. The weights of the nodes were drawn uniformly from the interval $[1,10]$. Manipulation of the quantum pulses is done using the Pulser [63] Python packages. Emulations of $N\leq 12$ qubits are done using the QuTip backend [64], while those with $N>12$ are done using emu-mps [65], an open-access tool for matrix-product state (MPS) simulations of neutral-atoms quantum systems. The emulator’s parameters are kept in their default settings.
@@ V.1 Synthetic instances
At each iteration of the CG scheme, there are $|V|$ PSPs that must be solved over subgraphs $G_{v}$ for all $v\in V$. To ensure a control over the hardness of our instances, we make sure that all $G_{v}$ remain hard (i.e. non-trivial). We control the number of vertices $\lvert K_{v}\rvert$ (i.e. the number of compatible tours per vehicle class), the number of edges $\lvert E_{v}\rvert$ via the edge’s probability $p=\lvert E_{v}\rvert/\binom{\lvert K_{v}\rvert}{2}$ (i.e. the number of conflicting tours). We also control the number of vehicle classes $\lvert V\rvert$ (i.e. the number of independent PSPv to solve at each iteration). There are also parameters influencing the complexity of the RMP: the average number of allowed vehicle types per tours $\bar{N_{V}}=1/\lvert K\rvert\sum_{k=1}^{\lvert K\rvert}|V_{k}|$, from which we infer the total number of tours in the problem as $K=\lceil\bar{N_{V}}\lvert K\rvert\rceil$, the costs of each tour $C_{l}^{(T)}$ and vehicle types $C_{v}^{(v)}$, and the availability bounds $N_{v}^{\min}$ and $N_{v}^{\max}$ for each vehicle type $v$. Our synthetic instances have the following global parameters: $\lvert V\rvert=8$, graphs $G_{v}$ are Erdos-Renyi connected graphs with $p\sim 0.30$, $\bar{N_{V}}=2$, $C_{k}^{(T)}\in\mathcal{N}(10,5)$, $C_{v}^{(v)}\in\mathcal{N}(50,10)$, $N_{v}^{\min}=1$ and $N_{v}^{\max}$ where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and variance $\sigma^{2}$. The Instance classes are noted by the tuple $(|V|,|K_{v}|)$. We create 10 random instance (random generation of subgraphs and costs following the above distributions), for the $(8,8)$, $(8,10)$ and $(8,12)$ instances. For the $(8,14)$ instances, we generate 30 instances in order to perform a more in-depth analysis of the impact of the chosen solver’s performance on the CG scheme.
@@ VI Conclusion
We showed in this paper that novel register embedding and pulse design protocols for analog neutral-atoms quantum computers can enhance the performance of hybrid column generation schemes. We showed that these protocols can compete with other classical solvers for the pricing subproblems in returning a large set of high quality and diversity columns at each iteration. We also introduced a new greedy postprocessing method called Make_Diff which seeks to modify degenerate samples in order to obtain a set of potential columns that are all non-degenerate. Importantly, our quantum protocol with this postprocessing most closely mimics the behavior of an idealized solver that finds the $M$ best columns to add to the reduced master problem.
The focus of this paper was on the fleet assignment problem, which acted as a hard combinatorial problem on which to test our hypotheses concerning the design and performance of analog quantum primitives in column generations schemes. We note that the conclusions and the methods presented here can be applied to other practical settings. In machine job scheduling, for example, jobs must be assigned to machines in a way that respects machine capabilities and job durations, aiming to optimize various cost metrics. Likewise, in airline and public transit scheduling, aircraft, or buses must be assigned to flight or route sequences while accounting for various constraints. In such industrial settings, one may need to expand the workflow to include branching (as was done in Ref. [67]) or the generation of cuts [68].
These settings all represent prime places on which an industrial quantum advantage may be sought. To achieve such an advantage, our paper places a threshold to meet: the quantum protocol must be able to return a set of $M$ high quality and high diversity columns at each iterations faster than competing classical methods, such as simulated annealing and tensor networks. This remains an open question, although the experimental results of quantum annealing protocols to solve large-scale maximum independent set problems on neutral-atoms hardware [69, 62, 21] are encouraging.
@@ Acknowledgements.
We thank Wesley Coelho, Yassine Naghmouchi and Hossein Sadeghi for valuable discussions and support. This research was financially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and Pasqal Canada through an Alliance grant (ALLRP 590810 - 23). This work made use of compute resources by Calcul Québec and the Digital Research Alliance of Canada.